# DriveDreamer4D: World Models Are Effective Data Machines for 4D Driving Scene Representation

### v2

@@ Abstract
Closed-loop simulation is essential for advancing end-to-end autonomous driving systems. Contemporary sensor simulation methods, such as NeRF and 3DGS, rely predominantly on conditions closely aligned with training data distributions, which are largely confined to forward-driving scenarios. Consequently, these methods face limitations when rendering complex maneuvers (e.g., lane change, acceleration, deceleration). Recent advancements in autonomous-driving world models have demonstrated the potential to generate diverse driving videos. However, these approaches remain constrained to 2D video generation, inherently lacking the spatiotemporal coherence required to capture intricacies of dynamic driving environments. In this paper, we introduce DriveDreamer4D, which enhances 4D driving scene representation leveraging world model priors. Specifically, we utilize the world model as a data machine to synthesize novel trajectory videos based on real-world driving data. Notably, we explicitly leverage structured conditions to control the spatial-temporal consistency of foreground and background elements, thus the generated data adheres closely to traffic constraints. To our knowledge, DriveDreamer4D is the first to utilize video generation models for improving 4D reconstruction in driving scenarios. Experimental results reveal that DriveDreamer4D significantly enhances generation quality under novel trajectory views, achieving a relative improvement in FID by 24.5%, 39.0%, and 10.5% compared to PVG, $\text{S}^{3}$aussian, and Deformable-GS. Moreover, DriveDreamer4D markedly enhances the spatiotemporal coherence of driving agents, which is verified by a comprehensive user study and the relative increases of 20.3%, 42.0%, and 13.7% in the NTA-IoU metric.
@@ 1 Introduction
End-to-end planning [24, 25, 28], which directly maps sensor inputs to control signals, is among the most critical and promising tasks in autonomous driving. However, current open-loop evaluations are inadequate for accurately assessing end-to-end planning algorithms, highlighting an urgent need for enhanced evaluation methods [36, 75]. A compelling solution lies in closed-loop evaluations within real-world scenarios, which require retrieving sensor data from arbitrarily specified viewpoints. This necessitates constructing a 4D driving scene representation capable of reconstructing complex, dynamic driving environments.
@@
In this paper, we introduce DriveDreamer4D, which improves 4D driving scene representation by integrating priors from autonomous driving world models. Our approach utilizes an autonomous driving world model [57, 76] as a generative engine, synthesizing novel trajectory video data that densifies real-world driving datasets for enhanced training. Notably, we propose the Novel Trajectory Generation Module (NTGM) to generate diverse structured traffic conditions, and DriveDreamer4D applies these conditions to independently regulate the motion dynamics of foreground and background elements in complex driving environments. These conditions undergo view projection synchronized with vehicle maneuvers, ensuring that the synthesized data rigorously adheres to the spatiotemporal constraints of 4D driving scenarios. To the best of our knowledge, DriveDreamer4D is the first framework to harness video generation models for elevating 4D scene reconstruction quality in autonomous driving, providing richly varied viewpoint data for scenarios including lane change, acceleration, and deceleration. As shown in Fig. 1, experiment results demonstrate that DriveDreamer4D significantly enhances generation fidelity for novel trajectory viewpoints, achieving a relative improvement in FID by 24.5%, 39.0%, and 10.5% compared to PVG [8], $\text{S}^{3}$aussian [26], and Deformable-GS [69]. Besides, DriveDreamer4D fortifies the spatiotemporal coherence between foreground and background elements, with respective increases of 20.3%, 42.0%, and 13.7% in the NTA-IoU metric. Furthermore, a comprehensive user study confirms that the average win rate of DriveDreamer4D exceeds 80%, compared to three baselines.
The primary contributions of this work are as follows: (1) We present DriveDreamer4D, the first framework to leverage world model priors for advancing 4D scene reconstruction in autonomous driving. (2) The NTGM is proposed to automate the generation of diverse structured conditions, enabling DriveDreamer4D to produce novel trajectory videos with complex maneuvers. By explicitly incorporating structured conditions, DriveDreamer4D ensures spatial-temporal consistency across both foreground and background elements. (3) We perform comprehensive experiments to validate that DriveDreamer4D notably enhances generation quality across novel trajectory viewpoints, as well as the spatiotemporal coherence of driving scene elements.
@@ 2.1 Driving Scene Representation
NeRF and 3DGS have emerged as leading approaches for 3D scene representation. NeRF models [41, 2, 3, 42] continuous volumetric scenes using multi-layer perceptron (MLP) networks, enabling highly detailed scene reconstructions with remarkable rendering quality. More recently, 3DGS [30, 72] introduces an innovative method by defining a set of anisotropic Gaussians in 3D space, leveraging adaptive density control to achieve high-quality renderings from sparse point cloud inputs. Several works have extended NeRF [68, 66, 27, 39, 48, 54, 16] or 3DGS [65, 77, 8, 73, 26, 10] to autonomous driving scenarios. Given the dynamic nature of driving environments, there has also been significant effort in modeling 4D driving scene representations. Some approaches encode time as an additional input to parameterize 4D scenes [1, 11, 35, 38, 44, 52, 26], while others represent scenes as a composition of moving object models alongside a static background model [68, 33, 43, 55, 61, 63]. Despite these advancements, methods based on NeRF and 3DGS face limitations tied to the density of input data. These techniques can only render scenes effectively when sensor data closely matches the training data distribution, which is typically confined to forward-driving scenarios.
@@ 2.2 World Models
The world model module predicts possible future world states as a function of imagined action sequences proposed by the actor [34, 78]. Approaches such as [58, 64, 74, 15, 17, 4, 20, 40, 21, 5, 32, 70, 22, 62] simulate environments through video generation controlled by free-text actions. At the forefront of this evolution is Sora [6], which leverages advanced generative techniques to produce intricate visual sequences that respect the fundamental laws of physics. This ability to deeply understand and simulate the environment not only improves video generation quality but also has substantial implications for real-world driving scenarios. Autonomous driving world models [57, 76, 59, 23, 14, 67] employ predictive methodologies to interpret driving environments, thereby generating realistic driving scenarios and learning key driving elements and policies from video data. Although these models successfully produce diverse driving video data conditioned on complex driving actions, they remain limited to 2D outputs and lack the spatial-temporal coherence needed to accurately capture the complexities of dynamic driving environments.
@@ 2.3 Diffusion Prior for 3D Representation
Figure 2: The overall framework of DriveDreamer4D. Initially, by altering the actions of the original trajectory (e.g., steering angle, speed), new trajectories can be obtained. Conditioned on the first frame and the structured information (3D bounding boxes, HDMap) from the new trajectory, the novel trajectory videos are generated. Subsequently, both the original and novel trajectory videos are mixed to train the 4D Gaussian Splatting model.
@@ 3.1 Preliminary
3.1.2 World Models for Controllable Driving Video Generation
@@ 3.2 DriveDreamer4D
The overall pipeline of DriveDreamer4D is depicted in Fig. 2. In the upper part, the Novel Trajectory Generation Module (NTGM) is proposed to adjust original trajectory actions such as steering angle and speed to generate new trajectories. These novel trajectories provide fresh perspectives for extracting structured information like 3D boxes and HDMap details. Subsequently, a controllable video diffusion model synthesizes videos from these updated viewpoints, incorporating specific priors associated with the modified trajectories. In the lower part, both original and novel trajectory videos are integrated to optimize the 4DGS model. In the following sections, we delve into the details of novel trajectories video generation and then introduce 4D reconstruction with video diffusion priors.
@@ 3.2.1 Novel Trajectory Video Generation
As previously mentioned, traditional 4DGS methods are limited in rendering complex maneuvers, largely due to the training data being dominated by straightforward driving scenarios. To overcome this, DriveDreamer4D leverages world models priors to generate diverse viewpoint data, enhancing the 4D scene representation. To achieve this, we propose the NTGM, which is designed to create new trajectories that serve as input for the world model, enabling the automated generation of complex maneuver data. NTGM comprises two main components: (1) novel trajectory proposal, (2) trajectory safety assessment. In the novel trajectory proposal stage, text-to-trajectory [76] can be adopted to automatically generate diverse complex trajectories. Additionally, trajectories can be custom-designed to meet specific requirements, allowing for tailored data generation based on precise needs. The overview of the custom-designed trajectory proposal (e.g., lane change) and trajectory safety assessment is shown in the Algo. 1. In a specific driving scenario, the original trajectory in the world coordinate system can be readily acquired as $\mathcal{T}_{\text{ori}}^{\text{world}}=\{p_{i}^{\text{world}}\}_{i=0}^{K}$, where $K$ denotes the number of frames and $p_{i}^{\text{world}}\in\mathbb{R}^{3}$ refers to the position of the ego-vehicle at the $i$-th frame. To propose novel trajectories, the original trajectory $\mathcal{T}_{\text{ori}}^{\text{world}}$ is transformed into the ego-vehicle coordinate system of the first frame, denoted as $\mathcal{T}_{\text{ori}}^{\text{EgoStart}}$ and computed as:
@@
3.2.2 4D Reconstruction with Video Diffusion Prior
Based on the video diffusion prior, we can generate novel videos with diverse trajectories, enhancing the 4D reconstruction capabilities across different baselines [8, 26, 69]. Specifically, to train the 4DGS with video diffusion prior, it is essential to construct a hybrid dataset $D_{hybrid}$, which combines the original trajectory dataset $D_{\text{ori}}$ with the novel trajectory dataset $D_{\text{novel}}$. The balance between these datasets can be adjusted through the hyper-parameter $\beta$, allowing us to control the 4DGS scene reconstruction performance for both original and novel trajectories. This relationship is formulated as $D_{hybrid}=(1-\beta)D_{\text{ori}}+\beta D_{\text{novel}}$. The loss function $\mathcal{L}_{\text{novel}}$ for optimizing the 4DGS with generated data is akin to [8, 26], defined as follows:
$\mathcal{L}_{\text{novel}}(\phi^{\prime})$ $=\lambda_{\text{RGB}}\|\hat{I}_{\text{novel}}-I_{\text{novel}}\|_{1}$ (8)
$\quad+\lambda_{\text{SSIM}}\textrm{SSIM}(\hat{I}_{\text{novel}},I_{\text{novel}}),$
where $I_{\text{novel}}$ represents the generated images corresponding to the novel trajectories as described in Sec. 3.2.1, and $\hat{I}_{\text{novel}}$ denotes the rendered images under the novel trajectories via differentiable splatting [71]. Notably, different from [8, 26], depth maps are not employed as constraints in the optimization of 4DGS when using the generated dataset $D_{\text{novel}}$. The limitation arises from the fact that LiDAR point cloud data is exclusively collected for the original trajectory. When these LiDAR points are projected onto a new trajectory, it cannot produce a complete depth map for the new perspective, as something visible in the novel trajectory may have been occluded in the original view. Consequently, the incorporation of such depth maps does not facilitate the optimization of the 4DGS model. The more details are described in Sec. 4.3. The overall loss function for mixed training is defined as follows:
$\mathcal{L}(\phi^{\prime})=\mathcal{L}_{\text{ori}}+\lambda_{\text{novel}}\mathcal{L}_{\text{novel}}.$ (9)
@@
DriveDreamer4D with $\text{S}^{3}$aussian 0.491 53.36 0.474 52.60 0.445 52.57 0.470 52.84
@@
DriveDreamer4D with Deformable-GS 0.322 52.90 0.370 52.50 0.404 53.79 0.365 53.06
@@ 4 Experiments
In this section, we first outline the experimental setup, detailing the dataset, implementation details, and evaluation metrics. We then provide both quantitative and qualitative evidence demonstrating that the proposed DriveDreamer4D significantly enhances generation quality for novel trajectory viewpoints and improves the spatiotemporal coherence of foreground and background components.
@@ 4.1 Experiment Setup
Implementation Details. To demonstrate the versatility and robustness of DriveDreamer4D, we incorporate various 4DGS baselines into our pipeline, including Deformable-GS [69], $\text{S}^{3}$aussian [26], and PVG [8]. For a fair comparison, LiDAR supervision is introduced to Deformable-GS. During training, scenes are segmented into multiple clips, each containing 40 frames, aligned with the generative model’s output length. We use only forward-facing camera data and standardize the resolution across methods to $640\times 960$. Our model is trained for 50,000 iterations using the Adam optimizer [31], following the learning rate schedule used for 3D Gaussian Splatting. Training strategies and hyperparameters align with each baseline’s original setup, with each model trained for 50,000 iterations.
@@
PVG 105.29
DriveDreamer4D with PVG 79.54
$\text{S}^{3}$aussian 124.90
DriveDreamer4D with $\text{S}^{3}$aussian 76.24
Deformable-GS 92.34
DriveDreamer4D with Deformable-GS 82.67
@@
Figure 3: Qualitative comparisons of novel trajectory renderings during speed change scenarios. The orange boxes highlight that DriveDreamer4D significantly enhances the rendering quality across various baseline methods (PVG [8], $\text{S}^{3}$aussian [26], Deformable-GS [69]).
Additionally, in lane change scenarios, we observe inaccuracies in relative positioning, as well as frequent occurrences of artifacts such as flying points and ghosting, which notably degrade image quality. To assess this, we employ the FID metric [19], which quantifies differences in feature distribution between rendered novel trajectory images and original trajectory images. This metric effectively reflects visual quality and is particularly sensitive to artifacts like flying points and ghosting, providing a robust measure of image fidelity in these complex scenes. Finally, we conduct a user study to evaluate the generation quality. Specifically, we compare the visual results of each baseline method with its DriveDreamer4D enhanced version across three distinct novel trajectories. The evaluation criteria focus on overall video quality, with particular attention to foreground objects like vehicles. For each comparison, participants were asked to select the option they found most favorable. Further details are provided in the supplement.
Figure 4: Qualitative comparisons of novel trajectory renderings during lane change scenarios. The orange boxes highlight that DriveDreamer4D significantly enhances the rendering quality across various baseline methods (PVG [8], $\text{S}^{3}$aussian [26], Deformable-GS [69]).
4.2 Comparison with Different 4DGS Baselines
Quantitative Results. As demonstrated in Tab. 1, integrating DriveDreamer4D with different 4DGS algorithms consistently yields superior NTA-IoU and NTL-IoU scores across diverse, complex maneuvers (e.g., lane changes, acceleration, and deceleration), significantly outperforming the baseline methods. Specifically, with DriveDreamer4D, the average NTL-IoU scores across three baselines (PVG [8], $\text{S}^{3}$aussian [26], Deformable-GS [69]) are relatively enhanced by 1.5%, 3.5%, and 1.4%, underscoring DriveDreamer4D’s capability to significantly improve the spatiotemporal coherence of background lanes. Moreover, rendering dynamic foreground agents in intricate driving scenarios presents substantial challenges. However, DriveDreamer4D facilitates a relative improvement in average NTA-IoU for these baselines by 20.3%, 42.0%, and 13.7%, thereby markedly enhancing the spatiotemporal coherence of foreground agents in 4D rendering of driving scenarios.
In addition to verifying the spatiotemporal consistency of rendered novel trajectory views, we leverage the FID metric to assess rendering quality under novel trajectories. Given that acceleration and deceleration scenarios yield rendered views with distributional similarities to ground truth, limiting FID’s discriminative capability across algorithms, our FID comparisons focus specifically on lane change scenarios. Experiment results, as presented in Tab. 2, indicate that our method substantially outperforms the baseline methods (PVG [8], $\text{S}^{3}$aussian [26], Deformable-GS [69]), with FID relative improvements of 24.5%, 39.0%, and 10.5%. These results highlight DriveDreamer4D’s capability to enhance generation quality for novel trajectory viewpoints.
Finally, we conducted a user study to evaluate the rendering quality of different methods on novel trajectories, with a specific focus on foreground agents. For each method, we generated three novel trajectory views—lane change, acceleration, and deceleration—across eight scenes from the Waymo dataset. Participants were then asked to select the renderings they found most visually favorable in each comparison. The DriveDreamer4D win rates from this study, shown in Tab. 3, reveal a significant user preference for our method’s renderings.
@@ 4.2 Comparison with Different 4DGS Baselines
PVG [8] 100.0% 85.7% 79.4% 88.6%
$\text{S}^{3}$aussian [26] 100.0% 96.9% 90.6% 95.8%
Deformable-GS [69] 93.8% 81.2% 65.6% 80.2%

### v3

@@ Abstract
Closed-loop simulation is essential for advancing end-to-end autonomous driving systems. Contemporary sensor simulation methods, such as NeRF and 3DGS, rely predominantly on conditions closely aligned with training data distributions, which are largely confined to forward-driving scenarios. Consequently, these methods face limitations when rendering complex maneuvers (e.g., lane change, acceleration, deceleration). Recent advancements in autonomous-driving world models have demonstrated the potential to generate diverse driving videos. However, these approaches remain constrained to 2D video generation, inherently lacking the spatiotemporal coherence required to capture intricacies of dynamic driving environments. In this paper, we introduce DriveDreamer4D, which enhances 4D driving scene representation leveraging world model priors. Specifically, we utilize the world model as a data machine to synthesize novel trajectory videos, where structured conditions are explicitly leveraged to control the spatial-temporal consistency of traffic elements. Besides, the cousin data training strategy is proposed to facilitate merging real and synthetic data for optimizing 4DGS. To our knowledge, DriveDreamer4D is the first to utilize video generation models for improving 4D reconstruction in driving scenarios. Experimental results reveal that DriveDreamer4D significantly enhances generation quality under novel trajectory views, achieving a relative improvement in FID by 32.1%, 46.4%, and 16.3% compared to PVG, $\text{S}^{3}$aussian, and Deformable-GS. Moreover, DriveDreamer4D markedly enhances the spatiotemporal coherence of driving agents, which is verified by a comprehensive user study and the relative increases of 22.6%, 43.5%, and 15.6% in the NTA-IoU metric.
@@ 1 Introduction
End-to-end planning [26, 27, 30], which directly maps sensor inputs to control signals, is among the most critical and promising tasks in autonomous driving. However, current open-loop evaluations are inadequate for accurately assessing end-to-end planning algorithms, highlighting an urgent need for enhanced evaluation methods [39, 80, 37]. A compelling solution lies in closed-loop evaluations within real-world scenarios, which require retrieving sensor data from arbitrarily specified viewpoints. This necessitates constructing a 4D driving scene representation capable of reconstructing complex, dynamic driving environments.
@@
In this paper, we introduce DriveDreamer4D, which improves 4D driving scene representation by integrating priors from autonomous driving world models. Our approach utilizes an autonomous driving world model [81] as a generative engine, synthesizing novel trajectory video data that densifies real-world driving datasets for enhanced training. Notably, we propose the Novel Trajectory Generation Module (NTGM) to generate diverse structured traffic conditions, and DriveDreamer4D applies these conditions to independently regulate the motion dynamics of foreground and background elements in complex driving environments. These conditions undergo view projection synchronized with vehicle maneuvers, ensuring that the synthesized data adheres to the spatiotemporal constraints. Subsequently, the Cousin Data Training Strategy (CDTS) is proposed to merge temporal-aligned real and synthetic data for training 4DGS. In CDTS, a regularization loss is further incorporated to ensure perceptual coherence. To the best of our knowledge, DriveDreamer4D is the first framework to harness video generation models for elevating 4D scene reconstruction quality in autonomous driving, providing richly varied viewpoint data for scenarios including lane change, acceleration, and deceleration. As shown in Fig. 1, experiment results demonstrate that DriveDreamer4D significantly enhances generation fidelity for novel trajectory viewpoints, achieving a relative improvement in FID by 32.1%, 46.4%, and 16.3% compared to PVG [9], $\text{S}^{3}$aussian [28], and Deformable-GS [74]. Besides, DriveDreamer4D fortifies the spatiotemporal coherence between foreground and background elements, with respective increases of 22.6%, 43.5%, and 15.6% in the NTA-IoU metric. Furthermore, a comprehensive user study confirms that the average win rate of DriveDreamer4D exceeds 80%, compared to three baselines.
The primary contributions of this work are as follows: (1) We present DriveDreamer4D, the first framework to leverage world model priors for advancing 4D scene reconstruction in autonomous driving. (2) The NTGM is proposed to automate the generation of structured conditions, allowing DriveDreamer4D to create novel trajectory videos with complex maneuvers while ensuring spatial-temporal consistency. Additionally, the CDTS is introduced to merge temporal-aligned real and synthetic data for training 4DGS, using a regularization loss to maintain perceptual coherence. (3) We perform comprehensive experiments to validate that DriveDreamer4D notably enhances generation quality across novel trajectory viewpoints, as well as the spatiotemporal coherence of driving scene elements.
@@ 2.1 Driving Scene Representation
NeRF and 3DGS have emerged as leading approaches for 3D scene representation. NeRF models [45, 2, 3, 46] continuous volumetric scenes using multi-layer perceptron (MLP) networks, enabling highly detailed scene reconstructions with remarkable rendering quality. More recently, 3DGS [32, 77] introduces an innovative method by defining a set of anisotropic Gaussians in 3D space, leveraging adaptive density control to achieve high-quality renderings from sparse point cloud inputs. Several works have extended NeRF [73, 71, 29, 43, 52, 58, 18, 12] or 3DGS [70, 82, 9, 78, 28, 11, 40] to autonomous driving scenarios. Given the dynamic nature of driving environments, there has also been significant effort in modeling 4D driving scene representations. Some approaches encode time as an additional input to parameterize 4D scenes [1, 13, 38, 42, 48, 56, 28], while others represent scenes as a composition of moving object models alongside a static background model [73, 35, 47, 59, 66, 68]. Despite these advancements, methods based on NeRF and 3DGS face limitations tied to the density of input data. These techniques can only render scenes effectively when sensor data closely matches the training data distribution, which is typically confined to forward-driving scenarios.
@@ 2.2 World Models
The world model module predicts possible future world states as a function of imagined action sequences proposed by the actor [36, 83]. Approaches such as [63, 69, 79, 17, 19, 4, 22, 44, 23, 5, 34, 75, 24, 67, 62] simulate environments through video generation controlled by free-text actions. At the forefront of this evolution is Sora [6], which leverages advanced generative techniques to produce intricate visual sequences that respect the fundamental laws of physics. This ability to deeply understand and simulate the environment not only improves video generation quality but also has substantial implications for real-world driving scenarios. Autonomous driving world models [61, 81, 64, 25, 16, 72] employ predictive methodologies to interpret driving environments, thereby generating realistic driving scenarios and learning key driving elements and policies from video data. Although these models successfully produce diverse driving video data conditioned on complex driving actions, they remain limited to 2D outputs and lack the spatial-temporal coherence needed to accurately capture the complexities of dynamic driving environments.
@@ 2.3 Diffusion Prior for 3D Representation
Figure 2: The overall framework of DriveDreamer4D. Initially, by altering the actions of the original trajectory (e.g., steering angle, speed), new trajectories can be obtained. Conditioned on the first frame and the structured information (3D bounding boxes, HDMap) from the new trajectory, the novel trajectory videos are generated. Subsequently, the temporal-aligned cousin pair (original and novel trajectory videos) are merged to optimize the 4D Gaussian Splatting model, where a regularization loss is calculated to ensure perceptual coherence.
@@ 3.1 Preliminary
3.1.2 World Models for Driving Video Generation
@@ 3.2 DriveDreamer4D
The overall pipeline of DriveDreamer4D is depicted in Fig. 2. In the upper part, the Novel Trajectory Generation Module (NTGM) is proposed to adjust driving actions (e.g., steering angle, speed) to generate new trajectories. These novel trajectories provide new perspectives for extracting structured information like 3D boxes and HDMap. Subsequently, a controllable video diffusion model synthesizes videos from these updated viewpoints, incorporating specific priors associated with the modified trajectories. In the lower part, the Cousin Data Training Strategy (CDTS) is introduced to combine the temporal-aligned original and generated data for optimizing the 4DGS model, where a regularization loss is calculated to impose perceptual coherence. In the following sections, we delve into the details of novel trajectories video generation and then introduce the CDTS for 4D reconstruction.
@@ 3.2.1 Novel Trajectory Video Generation
As previously mentioned, traditional 4DGS methods are limited in rendering complex maneuvers, largely due to the training data being dominated by straightforward driving scenarios. To overcome this, DriveDreamer4D leverages world model priors to generate diverse viewpoint data, enhancing the 4D scene representation. To achieve this, we propose the NTGM, which is designed to create new trajectories that serve as input for the world model, enabling the automated generation of complex maneuver data. NTGM comprises two main components: (1) novel trajectory proposal, (2) trajectory safety assessment. In the novel trajectory proposal stage, text-to-trajectory [81] can be adopted to automatically generate diverse complex trajectories. Additionally, trajectories can be custom-designed to meet specific requirements, allowing for tailored data generation based on precise needs. The overview of the custom-designed trajectory proposal (e.g., lane change) and trajectory safety assessment is shown in the Algo. 1. In a specific driving scenario, the original trajectory in the world coordinate system can be readily acquired as $\mathcal{T}_{\text{ori}}^{\text{world}}=\{p_{i}^{\text{world}}\}_{i=0}^{K}$, where $K$ denotes the number of frames and $p_{i}^{\text{world}}\in\mathbb{R}^{3}$ refers to the position of the ego-vehicle at the $i$-th frame. To propose novel trajectories, the original trajectory $\mathcal{T}_{\text{ori}}^{\text{world}}$ is transformed into the ego-vehicle coordinate system of the first frame, denoted as $\mathcal{T}_{\text{ori}}^{\text{EgoStart}}$ and computed as:
@@
DriveDreamer4D with $\text{S}^{3}$aussian 0.495 53.42 0.484 52.63 0.445 52.69 0.475 52.91
@@
DriveDreamer4D with Deformable-GS 0.335 52.93 0.371 52.77 0.406 53.79 0.371 53.16
@@
3.2.2 Cousin Data Training Strategy
To better integrate generated data for training 4DGS, we propose the CDTS. Specifically, we construct temporally aligned cousin pair data as a minimal training batch:
$\text{BatchStack}(\{\hat{I}_{\text{ori},t}\}_{t=0}^{T},\{\hat{I}_{\text{novel},t}\}_{t=0}^{T}),$ (8)
where $\text{BatchStack}(\cdot)$ is the data processor to stack temporal-aligned $\{\hat{I}_{\text{ori},t}\}_{t=0}^{T}$ and $\{\hat{I}_{\text{novel},t}\}_{t=0}^{T}$ into a training batch. By leveraging real and synthetic data aligned at each timestep, CDTS mitigates the data gap in 4DGS training, enhancing the model’s ability to learn consistent representations across real and synthetic data. To optimize 4DGS, the temporal-aligned cousin pair is input per step before gradient optimization. The loss function $\mathcal{L}_{\text{ori}}$ for the original data is defined in Eq. 4. And the loss function $\mathcal{L}_{\text{novel}}$ for the generated data is akin to [9, 28], defined as follows:
$\mathcal{L}_{\text{novel}}(\phi^{\prime})$ $=\lambda_{\text{1}}\|\hat{I}_{\text{novel}}-I_{\text{novel}}\|_{1}$ (9)
$\quad+\lambda_{\text{3}}\textrm{SSIM}(\hat{I}_{\text{novel}},I_{\text{novel}}),$
where $I_{\text{novel}}$ represents the generated images corresponding to the novel trajectories as described in Sec. 3.2.1, and $\hat{I}_{\text{novel}}$ denotes the rendered images under the novel trajectories via differentiable splatting [76]. Notably, different from [9, 28], depth maps are not employed as constraints in the optimization of 4DGS when using the generated dataset $D_{\text{novel}}$. The limitation arises from the fact that LiDAR point cloud data is exclusively collected for the original trajectory. When these LiDAR points are projected onto a new trajectory, it cannot produce a complete depth map for the new perspective, as something visible in the novel trajectory may have been occluded in the original view. Consequently, the incorporation of such depth maps does not facilitate the optimization of the 4DGS model. More details are described in Sec. 4.3. Additionally, we propose a regularization loss to enhance the perceptual coherence, defined as follows:
$\mathcal{L}_{\text{reg}}(\phi^{\prime})=\|\mathcal{F}_{p}(\hat{I}_{\text{ori}})-\mathcal{F}_{p}(\hat{I}_{\text{novel}})\|_{1},$ (10)
where $\mathcal{F}_{p}(\cdot)$ denotes the perception feature extraction model [21]. The overall loss function for mixed training is defined as follows:
$\mathcal{L}(\phi^{\prime})=\mathcal{L}_{\text{ori}}+\lambda_{\text{novel}}\mathcal{L}_{\text{novel}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}.$ (11)
@@ 4 Experiments
In this section, we first outline the experimental setup, including the dataset, implementation details, and evaluation metrics. Subsequently, both quantitative and qualitative evidence are provided to demonstrate that the proposed DriveDreamer4D significantly enhances rendering quality for novel trajectory viewpoints and improves the spatiotemporal coherence of foreground and background components. Finally, we conduct an ablation study on the hyperparameter settings, as well as the effects of depth loss and the proposed CDTS, including temporal-aligned pairs, and regularization loss.
@@ 4.1 Experiment Setup
Implementation Details. To demonstrate the versatility and robustness of DriveDreamer4D, we incorporate various 4DGS baselines into our pipeline, including Deformable-GS [74], $\text{S}^{3}$aussian [28], and PVG [9]. For a fair comparison, LiDAR supervision is introduced to Deformable-GS. During training, scenes are segmented into multiple clips, each containing 40 frames, aligned with the generative model’s output length. We use only forward-facing camera data and standardize the resolution across methods to $640\times 960$. Our models are trained using the Adam optimizer [33], following the learning rate schedule used for 3D Gaussian Splatting [32]. Hyperparameter settings are aligned with each baseline [74, 28, 9], and the training strategy remains the same, with the exception of the incorporation of CDTS.
Figure 3: Qualitative comparisons of novel trajectory renderings during lane change scenarios. The orange boxes highlight that DriveDreamer4D significantly enhances the rendering quality across various baselines (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]).
@@
PVG[9] 105.29
DriveDreamer4D with PVG 71.52
$\text{S}^{3}$aussian[28] 124.90
DriveDreamer4D with $\text{S}^{3}$aussian 66.93
Deformable-GS[74] 92.34
DriveDreamer4D with Deformable-GS 77.32
@@
Additionally, in lane change scenarios, we observe inaccuracies in relative positioning, as well as frequent occurrences of artifacts such as flying points and ghosting, which notably degrade image quality. To assess this, we employ the FID metric [21], which quantifies differences in feature distribution between rendered novel trajectory images and original trajectory images. This metric effectively reflects visual quality and is particularly sensitive to artifacts like flying points and ghosting, providing a robust measure of image fidelity in these complex scenes. Finally, a user study is conducted to evaluate the quality of the renderings, where participants compare the rendering results of each baseline with its DriveDreamer4D enhanced version across three novel trajectories. The evaluation criteria focus on overall video quality, with particular attention to foreground objects like vehicles. For each comparison, participants were asked to select the option they found most favorable. Further details are provided in the supplement.
@@
PVG [9] 100.0% 90.5% 89.1% 93.2%
$\text{S}^{3}$aussian [28] 100.0% 97.9% 92.2% 96.7%
Deformable-GS [74] 95.8% 83.5% 72.9% 84.1%